# A Fourier-Jacobi Dirichlet series attached to modular forms of $\textup{SO}(2,4)$

### v1

@@ Abstract.
We consider a Dirichlet series $D_{F,G}(s)$ attached to two modular forms $F$ and $G$ of an orthogonal group of real signature $(2,4)$, involving their Fourier–Jacobi coefficients. When $F$ is a Hecke eigenform and $G$ a Poincaré series, our main result gives that $D_{F,G}(s)$ is equal to the standard $L$-function attached to $F$, up to an explicit constant. To establish this, we use a correspondence between binary Hermitian forms and ideals of quaternion algebras, as established by Latimer, together with the fact that the even Clifford algebra of a three-dimensional definite quadratic space can be identified with a quaternion division algebra. Our work should be seen as a generalisation of a work of Kohnen and Skoruppa, whose result corresponds to the case of the orthogonal group of real signature $(2,3)$.
@@ 1. Introduction
Let $F$ and $G$ be two degree two Siegel cusp forms of even weight $k\geq 0$. In their groundbreaking work [KS89], Kohnen and Skoruppa define and study the Dirichlet series
@@
where $\phi_{N}$ and $\psi_{N}$ are the $N$th Fourier–Jacobi coefficients of $F$ and $G$ respectively, and $\langle\,$$\rangle$ is the Petersson inner product on Jacobi forms of weight $k$ and index $N$. The significance of their work derives mainly from the fact that if one further assumes that $F$ is a Hecke eigenform and $G$ is in the Maass space, then this Dirichlet series is proportional to $Z_{F}(s)$, the spinor $L$-function attached to $F$. Actually, as explained in their paper, it is sufficient to establish such a relation for the case where $G$ is taken to be the lift of a Poincaré series, as such lifts generate the Maass space in the symplectic setting. In particular, they show in [KS89, Theorem 3],
$D_{F,\mathcal{P}_{k,D}}(s)=A(Q)Z_{F}(s),$
@@
where $Q$ denotes any quadratic form of discriminant $D$ representing $1$, and $A(Q)$ is the $Q$th Fourier coefficient of $F$. Here, $\mathcal{P}_{k,D}$ is a lift of a classical Poincare series associated to $k$ and $D$ (see [KS89, page 552]).
@@
In this paper, we obtain a result analogous to Kohnen and Skoruppa’s for modular forms of orthogonal groups of signature (2,4). We will first describe our result and then explain how it fits into a more general picture, alongside Kohnen and Skoruppa’s work.
Let $S$ denote an even, positive definite symmetric matrix of rank $2$, such that $\mathbb{Z}^{2}$ is a maximal lattice with respect to $S$. We put $K:=\mathbb{Q}\left(\sqrt{-\det S}\right)$. Let
$S_{0}:=\begin{pmatrix}&&1\\ &-S&\\ 1&&\end{pmatrix},\,\,\,S_{1}=\begin{pmatrix}&&1\\ &-S_{0}&\\ 1&&\end{pmatrix}$,
and write $(V_{i},\phi_{i})$ for the quadratic spaces associated to $S_{i}$, for $i=0,1$. We also write $\textup{SO}^{\phi_{i}}(V_{i})$ for the corresponding special orthogonal groups. Let now $F$ denote an orthogonal cusp form of integer even weight $k\geq 0$ with respect to a congruence subgroup $\Gamma_{S}$ of $\textup{SO}^{\phi_{1}}(V_{1})$. Assume $\{\phi_{m}\}_{m=1}^{\infty}$ are its Fourier–Jacobi coefficients. We also consider a Maass lift of a Poincaré series, given by
$\mathcal{P}_{k,D,r}(\tau^{\prime},z,\tau):=\sum_{N\geq 1}(V_{N}P_{k,D,r})(\tau,z)e(N\tau^{\prime}),$
where $V_{N}$ is an index-raising operator acting on Fourier–Jacobi forms of lattice index and $(D,r)\in\mathbb{Z}\times\mathbb{Z}^{2}$ is in the support of the maximal lattice $\mathbb{Z}^{2}$ with respect to $S$. Here, $(\tau^{\prime},z,\tau)$ is in the associated symmetric space (see 5). For any finite set of primes $\mathcal{P}$, we then consider the Dirichlet series
@@
$\mathcal{D}_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\,\forall p\in\mathcal{P}\end{subarray}}^{\infty}\langle\phi_{N},V_{N}P_{k,D,r}\rangle N^{-s},$
and set
$D_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s):=\zeta_{K,\mathcal{P}}(s-k+3)\zeta_{\mathcal{P}}(2s-2k+4)\mathcal{D}_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s),$
@@
where $\langle\ ,\ \rangle$ denotes the inner product on the space of Fourier–Jacobi forms of weight $k$ and (lattice) index $NS$, and $\zeta_{K}$ denotes the Dedekind zeta function attached to $K$.
We now take $D=-\ell q$, with $\ell\in\mathbb{Z}_{>0}$ and $q$ the level of $S$. We let $L_{0}:=\mathbb{Z}^{4}$ and then choose $\xi=(1,0,0,\ell)\in V_{0}$, for which $\phi_{0}[\xi]=\ell$ and $2\phi_{0}(\xi,L_{0})=\mathbb{Z}$. The main Theorem of this paper is the following.
@@
Assume $K$ has class number $1$ and it is not $\mathbb{Q}(\sqrt{-2})$. Choose $\ell$ so that it satisfies the conditions of Proposition 7.1. Assume $F$ is a Hecke eigenform for the corresponding Hecke algebra. Then, outside a finite set of primes $\mathcal{P}$, we have
$D_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)=A(\xi)L_{\mathcal{P}}(F;s-k+2),$
@@
where $L(F;s)$ is the standard $L$-function attached to $F$, and $A(\xi)$ is the $\xi^{th}$ Fourier coefficient of $F$. In particular, if $A(\xi)\neq 0$, then $D_{F,\mathcal{P}_{k,D,r},\mathcal{P}}(s)$ is proportional to $L_{\mathcal{P}}(F;s-k+2)$.
We mention here that we also have computational evidence which suggests that Proposition 7.1 can be extended to also include the case of $K=\mathbb{Q}(\sqrt{-2})$. However, we have not yet been able to prove it, and we hope to return to this in the near future.
Let us now explain how the above results can be seen in a more general setting and how this paper is related to the recent work [Psy25] of one of the authors. It is a well-known fact that there is an accidental isogeny of the symplectic group $\textup{Sp}_{2}(\mathbb{R})$ and of $\textup{SO}(2,3)$, the special orthogonal group of signature $(2,3)$. Moreover, the spinor $L$-function attached to $F$ above can be identified with the standard $L$-function attached to a holomorphic modular form on an appropriate orthogonal group of real signature $(2,3)$, as for example is discussed in [Shi04, Section 25]. In fact, there are more accidental isogenies between classical and orthogonal groups. Examples are $\textup{U}(2,2)$ and $\textup{SO}(2,4)$, and $\textup{Sp}(2,\mathbb{H})$ and $\textup{SO}(2,6)$, where $\mathbb{H}$ denotes the skew field of real quaternions. Modular forms can be defined for orthogonal groups of any real signature $(2,n+2)$, $n\geq 1$. Crucially, the associated symmetric space is of tube type and has the structure of a Shimura variety. It is therefore natural to ask whether the method of Kohnen and Skoruppa can be generalised to the orthogonal setting. This has been the subject of study in a recent work of one of the authors in [Psy25].
Indeed, consider the same setting as above, but with $S$ now having rank $n\geq 1$. We denote by $H(\xi)$ the subgroup of $\textup{SO}^{\phi_{0}}(V_{0})$ given by
$H(\xi)_{\mathbb{Q}}:=\{g\in\textup{SO}^{\phi_{0}}(V_{0})\mid g\xi=\xi\},$
which is a negative definite orthogonal group of rank $n+1$. It is one of the main results by Shimura in [Shi04] that the congruence subgroup of $\textup{SO}^{\phi_{0}}(V_{0})$ which fixes the lattice $L_{0}:=\mathbb{Z}^{n+2}$ acts on the set
$\left\{x\in V_{0}\mid\phi_{0}(x,x)=-D/q\textup{ and }2\phi_{0}(x,L_{0})=\mathbb{Z}\right\},$
and there are finitely many equivalence classes under this action. Let $\{\xi_{i}\}_{i=1}^{h}$ be representatives for this action. These representatives correspond to elements of the finite set
$H(\xi)_{\mathbb{Q}}\backslash H(\xi)_{\mathbb{A}}/(H(\xi)_{\mathbb{A}}\cap C),$ (1)
where $C=\{x\in\textup{SO}^{\phi_{0}}(V_{0})_{\mathbb{A}}\mid xL_{0}=L_{0}\}$. Assume $\{u_{i}\}_{i=1}^{h}$ are these elements. This is the analogue of the main theorem for binary quadratic forms of fixed discriminant.
Let also $\{f_{j}\}_{j=1}^{h}$ denote an orthonormal basis of simultaneous eigenforms on the set (1). Denote by $L(-,s)$ the standard $L$-function attached to either $F$ or any of the $f_{j}^{\prime}s$. We then have the following Theorem.
Theorem 1.2.
([Psy25, Theorem 8.2]) Outside a finite set of primes $\mathcal{P}$, $\mathcal{D}_{F,\mathcal{P}_{k,D,r}}(s)$ can be written as
$L_{\mathcal{P}}\left(F;s-k+(n+2)/2\right)\sum_{j=1}^{h}A_{f_{j}}L_{\mathcal{P}}\left(\overline{f_{j}};s-k+(n+3)/2\right)^{-1}\times\\ \times\sum_{i=1}^{h}\zeta_{\xi_{i},\mathcal{P}}(s-k+n+1)f_{j}(u_{i})\times\begin{cases}1&\textup{if }n\textup{ odd}\\ \zeta_{\mathcal{P}}(2s-2k+n+2)^{-1}&\textup{if }n\textup{ even}\end{cases}.$
Here, $\displaystyle{\zeta_{\xi_{i}}(s)=\sum_{d=1}^{\infty}n(\xi_{i};d)d^{-s}}$, with
$n(\xi_{i};d):=\#\left\{s\in\mathbb{Z}^{n}/dS\mathbb{Z}^{n}\mid D\equiv\frac{1}{2}qS^{-1}[s]\pmod{qd},\,\,\left(\frac{\frac{1}{2}qS^{-1}[s]-D}{qd},S^{-1}s,d\right)^{t}=\gamma\xi_{i},\right.\\ \left.\gamma\in\Gamma(L_{0})\right\}.$
and the $A_{f_{j}}$ are expressions depending on $f_{j}$ and the Fourier coefficients of $F$. For any zeta function, the subscript $\mathcal{P}$ means that we do not take into account the terms sharing factors with elements of $\mathcal{P}$.
It is therefore clear that to obtain a clear-cut Euler product, it is necessary to relate
$\sum_{i=1}^{h}\zeta_{\xi_{i},\mathcal{P}}(s-k+n+1)f_{j}(u_{i}),$
to the standard $L$-function attached to $\overline{f_{j}}$ for each $j=1,\cdots,h$. In [Psy25, Section 9], we restricted ourselves to orthogonal groups for which $h=1$ and obtained clear-cut Euler products in [Psy25, Theorem 9.2, Theorem 9.3].
@@
The main contribution of the present paper is to drop the condition $h=1$ for the $n=2$ case. Our strategy is to exploit the isogeny between the unitary group $\textup{U}(2,2)$ and $\textup{SO}(2,4)$. In particular, we identify the $\xi_{i}$’s with classes of binary Hermitian forms of discriminant $\ell$. According to Latimer in [Lat35], these correspond to specific classes of ideals in a certain quaternion algebra $B$, which allow us to identify each $\zeta_{\xi_{i}}(s)$ as the partial zeta function associated with classes of ideals in $B$. We then identify such a quaternion algebra as an even Clifford algebra associated to $\textup{SO}(3)$, and interpret the $f_{j}$’s as lifts to modular forms on the corresponding Clifford group. Our result then follows using Shimura’s results in [Shi04, Section 22.11].
Finally, we should mention here that Gritsenko in [Gri92] has considered the analogous problem for the case of Hermitian modular forms of degree two over the imaginary quadratic field $K=\mathbb{Q}(i)$. Gritsenko’s approach is to use factorisation methods in parabolic Hecke rings, and, in fact, is quite different from the techniques used in this paper. Indeed, our work is closer to the approach of Kohnen and Skoruppa [KS89], where the representation numbers of binary Hermitian forms in our work play an analogue role to the representation numbers of binary quadratic forms in their work. In fact, we also establish a relation of the standard $L$-function of $\textup{SO}(3)$ with the zeta functions $\zeta_{\xi_{i}}$’s (Theorem 10.5). These zeta functions count solutions of congruences and hence can potentially be used to compute the $L$-function explicitly. We do believe that our main result is compatible with Grisenko’s main Theorem in [Gri92, (5.3)], but we have not yet checked all the details. We hope to return to this in the near future.

### v2

@@ Abstract.
We consider a Dirichlet series $D(\bm{F},\bm{G};s)$ attached to two automorphic forms $\bm{F}$ and $\bm{G}$ of an orthogonal group of real signature $(2,4)$, involving their Fourier–Jacobi coefficients. When $\bm{F}$ is a Hecke eigenform and $\bm{G}$ a lift of a Jacobi-Poincaré series, our main result gives that $D(\bm{F},\bm{G};s)$ is equal to the standard $L$-function attached to $\bm{F}$, up to an explicit constant. To establish this, we use a correspondence between binary Hermitian forms and ideals of quaternion algebras, as established by Latimer, together with the fact that the even Clifford algebra of a three-dimensional definite quadratic space can be identified with a quaternion division algebra. Our work should be seen as a generalisation of a work of Kohnen and Skoruppa, whose result corresponds to the case of the orthogonal group of real signature $(2,3)$.
@@ 1. Introduction
Let $F$ and $G$ be two degree two Siegel cusp forms of even weight $k\geq 0$. In their groundbreaking work [kohnen_skoruppa], Kohnen and Skoruppa define and study the Dirichlet series
@@
where $\phi_{N}$ and $\psi_{N}$ are the $N$th Fourier–Jacobi coefficients of $F$ and $G$ respectively, and $\langle\,$$\rangle$ is the Petersson inner product on Jacobi forms of weight $k$ and index $N$. The significance of their work derives mainly from the fact that if one further assumes that $F$ is a Hecke eigenform and $G$ is in the Maass space, then this Dirichlet series is proportional to $Z_{F}(s)$, the spinor $L$-function attached to $F$. Actually, as explained in their paper, it is sufficient to establish such a relation for the case where $G$ is taken to be the lift of a Jacobi-Poincaré series, as such lifts generate the Maass space in the symplectic setting. In particular, they show in [kohnen_skoruppa, Theorem 3],
(1.1) $D_{F,\mathcal{P}_{k,D}}(s)=A(Q)Z_{F}(s),$
@@
where $Q$ denotes any quadratic form of discriminant $D$ representing $1$, and $A(Q)$ is the $Q$th Fourier coefficient of $F$. Here, $\mathcal{P}_{k,D}$ is a lift of a classical Poincare series associated to $k$ and $D$ (see [kohnen_skoruppa, page 552]).
@@
In this paper, we obtain a result analogous to Kohnen and Skoruppa’s for modular forms of orthogonal groups of signature $(2,4)$. We will first describe our result and then explain how it fits within a broader framework, alongside Kohnen and Skoruppa’s work.
Let $V$ denote a two-dimensional quadratic space and let $\phi:V\times V\longrightarrow\mathbb{Q}$ denote a positive definite, symmetric, quadratic form. We assume that there is some $v\in V$ such that $\phi(v,v)=1$ (but see also Remark 4.1). We then have that $(V,\phi)$ is isomorphic to $(K,N_{K/\mathbb{Q}})$, where $K$ is an imaginary quadratic field. For a fractional ideal $\mathfrak{m}$ of $K$, we write $L_{\mathfrak{m}}$ for the lattice in $V$ corresponding to the fractional ideal $\overline{\mathfrak{m}}\,\mathfrak{m}^{-1}$. It is known that $L_{\mathfrak{m}}$ is a $\mathbb{Z}$-maximal lattice in $(V,\phi)$ and all $\mathbb{Z}$-maximal lattices of $(V,\phi)$ are of this form. We set $L:=L_{\mathcal{O}_{K}}$, and after fixing a $\mathbb{Z}$-basis for $L$ we may write $L=\mathbb{Z}^{2}$, $V=\mathbb{Q}^{2}$ and $S$ for a positive definite matrix such that
$\phi(x,y)=\frac{1}{2}x^{t}Sy,$
for all $x,y\in\mathbb{Q}^{2}$.
Consider now the quadratic spaces $(V_{0},\phi_{0})$ and $(V_{1},\phi_{1})$, which are represented by the matrices
$S_{0}:=\begin{pmatrix}&&1\\ &-S&\\ 1&&\end{pmatrix},\,\,\,S_{1}:=\begin{pmatrix}&&1\\ &-S_{0}&\\ 1&&\end{pmatrix}$,
with respect to the $\mathbb{Z}$-maximal lattices $L_{0}:=\mathbb{Z}^{4}$ and $L_{1}:=\mathbb{Z}^{6}$, respectively. We denote by $G_{\mathbb{Q}}^{*}$ and $G_{\mathbb{Q}}$ the special orthogonal groups associated with $S_{0}$ and $S_{1}$, respectively.
Let now $F$ denote an orthogonal cusp form of integer even weight $k\geq 0$ with respect to the subgroup $\Gamma(L_{1}):=\{\gamma\in G_{\mathbb{Q}}\mid\gamma L_{1}=L_{1}\}$. If now $D:=\{g\in G_{\mathbb{A}}\mid gL_{1}=L_{1}\}$, it is well-known that the class number of $G_{\mathbb{Q}}$ with respect to $D$ is one, and therefore $F$ corresponds uniquely to an automorphic cusp form of weight $k$ on $G_{\mathbb{A}}$, which we denote by $\bm{F}$.
Now, for each fractional ideal $\mathfrak{m}$ of $K$, let $L_{\mathfrak{m},1}:=\mathbb{Z}^{2}+L_{\mathfrak{m}}+\mathbb{Z}^{2}$ denote the corresponding lattices in $V_{1}$. These are all $\mathbb{Z}$-maximal lattices in $V_{1}$, and because of the class number one property, we can pick isometries $\widetilde{h}_{\mathfrak{m}}\in G_{\mathbb{Q}}$ such that $L_{\mathfrak{m},1}=\widetilde{h}_{\mathfrak{m}}L_{1}$. After picking these isometries in a particular way (see Lemma 4.2), we use $\bm{F}$ to define modular forms $F_{\mathfrak{m}}:=\bm{F}(\widetilde{h}_{\mathfrak{m}};Z)$, which are then orthogonal modular forms with respect to $\Gamma(L_{\mathfrak{m},1})$. Moreover, each $F_{\mathfrak{m}}$ admits a Fourier-Jacobi expansion, and we write $\{\phi_{\mathfrak{m},N}\}_{N=1}^{\infty}$ for these coefficients. In particular, they are Jacobi cusp forms of weight $k$ and lattice index $(L_{\mathfrak{m}},N\phi)$.
Furthermore, for each $\mathfrak{m}$ and $D_{\mathfrak{m}}\in\mathbb{Z}_{<0}$, we can define a Jacobi-Poincaré series $P_{\mathfrak{m}}=P_{\mathfrak{m}}(k,D)$ of weight $k$ and lattice index $L_{\mathfrak{m}}$. We can then consider a lift of such Jacobi-Poincaré series, given by
$\mathcal{P}_{\mathfrak{m}}(\tau^{\prime},z,\tau):=\sum_{N\geq 1}(V_{N}P_{\mathfrak{m}})(\tau,z)e(N\tau^{\prime}),$
where $V_{N}$ is an index-raising operator acting on Jacobi forms of lattice index. This is, in fact, an element of the corresponding Maass space. Moreover, it can be shown that each such $\mathcal{P}_{\mathfrak{m}}$ arises as $\bm{P}(\widetilde{h}_{\mathfrak{m}};Z)$, where $\bm{P}$ is the unique automorphic form corresponding to $\mathcal{P}:=\mathcal{P}_{\mathcal{O}_{K}}$.
For any finite set of primes $\mathcal{Q}$, we then consider the Dirichlet series
@@
$D_{\mathcal{Q}}(F_{\mathfrak{m}},\mathcal{P}_{\mathfrak{m}};s):=\sum_{\begin{subarray}{c}N=1\\ (N,p)=1\,\forall p\in\mathcal{Q}\end{subarray}}^{\infty}\langle\phi_{\mathfrak{m},N},V_{N}P_{\mathfrak{m}}\rangle N^{-s},$
@@
where $\langle\ ,\ \rangle$ denotes the inner product on the space of Fourier–Jacobi forms of weight $k$ and (lattice) index $(L_{\mathfrak{m}},N\phi)$. Moreover, after picking a set of representatives for $\textup{Cl}(K)$, we set
$D_{\mathcal{Q}}(\bm{F},\bm{P};s):=\sum_{\mathfrak{m}\in\textup{Cl}(K)}\zeta_{\mathfrak{m},\mathcal{Q}}(s-k+3)D_{\mathcal{Q}}(F_{\mathfrak{m}},P_{\mathfrak{m}};s),$
where $\zeta_{\mathfrak{m}}(s)$ denotes the zeta function attached to the ideal class $\mathfrak{m}$ (see (6.2)). The main Theorem of this paper is the following.
@@
Assume $K=\mathbb{Q}(\sqrt{-m})$ with $m$ square-free and $m\not\equiv 2\pmod{4}$. Choose $\ell\in\mathbb{Z}_{\geq 1}$ so that it satisfies the conditions of Proposition 9.1, and set $\xi:=(1,0,0,\ell)$. Assume $F$ is an eigenform for the corresponding Hecke algebra. Then, for a finite set of primes $\mathcal{Q}$, we have
$\zeta_{\mathcal{Q}}(2s-2k+4)D_{\mathcal{Q}}(\bm{F},\bm{P};s)=A(\xi)L_{\mathcal{Q}}(F;s-k+2),$
@@
where $L(F;s)$ is the standard $L$-function attached to $F$, and $A(\xi)$ is the $\xi^{th}$ Fourier coefficient of $F$. In particular, if $A(\xi)\neq 0$, then $D_{\mathcal{Q}}(\bm{F},\bm{P};s)$ is proportional to $L_{\mathcal{Q}}(F;s-k+2)$.
We refer to Corollary 12.8 for a statement for the series $D(\bm{F},\bm{G};s)$ when $\bm{G}$ belongs to the space spanned by the forms $\bm{P}$ satisfying the conditions of the Theorem above. Here, $D(\bm{F},\bm{G};s)$ is defined similar to $D(\bm{F},\bm{P};s)$ (see (6.1)).
We now discuss the main steps and the intuition behind our results. Our main goal has been to obtain an analogous result to that of Kohnen and Skoruppa. To that end, we consider a modular form $F$ for the full integral group $\Gamma(L_{1})$, as well as a modular form $\mathcal{P}$, which is a lift of a Jacobi-Poincaré series, and hence an element of the corresponding Maass space (in fact, such lifts generate the Maass space). As we mentioned above, in the case $n=2$, the quadratic space $(V,\phi)$ is identified with $(K,N_{K/\mathbb{Q}})$, where $K$ is a quadratic imaginary field. Contrary to the result of Kohnen and Skoruppa, though, considering just the Dirichlet series $D(F,\mathcal{P};s)$ is not enough to obtain an Euler product. The reason for this is that, in general, there is not a single class of $\mathbb{Z}$-maximal lattices in $(V,\phi)$ (in the $n=1$ case, this is always true). It is, therefore, crucial to consider various Dirichlet series, parametrised by these classes. Denoting these classes by $L_{\mathfrak{m}}$, where $\mathfrak{m}$ runs over fractional ideals of $K$, our next step is to consider appropriate modular forms $F_{\mathfrak{m}}$ corresponding to $L_{\mathfrak{m}}$, which will be used to define these Dirichlet series. Because of the class number one property of the space $(V_{1},\phi_{1})$, $F$ can be identified with a unique automorphic form $\bm{F}$, which then can be used to produce the various $F_{\mathfrak{m}}$. These are modular forms for the congruence subgroups of fixing the lattices $L_{\mathfrak{m},1}:=\mathbb{Z}^{2}+L_{\mathfrak{m}}+\mathbb{Z}^{2}$ in $V_{1}$. By then considering the corresponding lifts $\mathcal{P}_{\mathfrak{m}}$ of the Jacobi-Poincaré series for the lattices $L_{\mathfrak{m}}$, we are able to define the Dirichlet series $D(F_{\mathfrak{m}},\mathcal{P}_{\mathfrak{m}};s)$ for all $\mathfrak{m}$.
The next key idea is that one can treat each $D(F_{\mathfrak{m}},\mathcal{P}_{\mathfrak{m}};s)$ in an almost uniform manner. For each $\mathfrak{m}$, we consider the set of vectors $\xi$ in $V_{0}$, which satisfy $\phi_{0}[\xi]=\ell$ and $2\phi_{0}(\xi,L_{\mathfrak{m},0})=\mathbb{Z}$ for some $\ell>0$. This set splits into a finite number of orbits under the action of $\Gamma(L_{\mathfrak{m},0})$, and we write $\xi_{\mathfrak{m},i}$ for a choice of representatives. Fixing $\xi:=(1,0,0,\ell)$, we let
$H(\xi):=\{g\in G_{\mathbb{Q}}^{*}\mid g\xi=\xi\}.$
This is a negative definite orthogonal group of rank $3$. Using Proposition 6.2, we can rewrite $D(F_{\mathfrak{m}},\mathcal{P}_{\mathfrak{m}};s)$ as an expression involving some zeta functions $\zeta(\xi_{\mathfrak{{m},i}};s)$ counting the number of solutions of certain congruences, and Dirichlet series $D(F_{\mathfrak{m}},\xi_{\mathfrak{m},i};s)$ involving the Fourier coefficients of $F$. Crucially, the $D(F_{\mathfrak{m}},\xi_{\mathfrak{m},i};s)$ are independent of $\mathfrak{m}$, and, thanks to work of Sugano [sugano], are related to the standard $L$-function attached to $F$, as well as to the standard $L$-functions of a basis of eigenforms $f_{j}$’s on $H(\xi)$ (see (8.7)). We note here that the work of Sugano plays in our paper the role that the classical work of Andrianov [andrianov] plays in the work of Kohnen and Skoruppa.
@@
Dealing with the $\zeta(\xi_{\mathfrak{{m},i}};s)$ is at the heart of this paper. Our strategy is to exploit the isogeny between the unitary group $\textup{U}(2,2)$ and $\textup{SO}(2,4)$. In particular, we identify the $\xi_{i}$’s with classes of binary Hermitian forms of discriminant $\ell$, and hence relate the series $\zeta(\xi_{\mathfrak{{m},i}};s)$ with $\mathfrak{m}$-primitive representation numbers of the Hermitian form represented by $\xi_{i}$. Latimer, in [Latimer1], showed that binary hermitian forms correspond to specific classes of ideals in a certain quaternion algebra $B$, which allow us to identify each $\zeta(\xi_{\mathfrak{{m},i}};s)$ as the partial zeta function associated with classes of ideals in $B$. We then identify such a quaternion algebra as an even Clifford algebra associated to $\textup{SO}(3)$, and interpret the $f_{j}$’s as lifts to modular forms on the corresponding Clifford group. The proof of our main result then follows using Shimura’s work in [shimura2004arithmetic, Section 22.11].
We now explain how the above results can be seen in a more general setting and how this paper is related to the recent work [psyroukis_orthogonal] of one of the authors. It is a well-known fact that there is an accidental isogeny of the symplectic group $\textup{Sp}_{2}(\mathbb{R})$ and of $\textup{SO}(2,3)$, the special orthogonal group of signature $(2,3)$. Moreover, the spinor $L$-function attached to $F$ above can be identified with the standard $L$-function attached to a holomorphic modular form on an appropriate orthogonal group of real signature $(2,3)$, as for example is discussed in [shimura2004arithmetic, Section 25]. In fact, there are more accidental isogenies between classical and orthogonal groups. Examples are $\textup{U}(2,2)$ and $\textup{SO}(2,4)$, and $\textup{Sp}(2,\mathbb{H})$ and $\textup{SO}(2,6)$, where $\mathbb{H}$ denotes the skew field of real quaternions. Modular forms can be defined for orthogonal groups of any real signature $(2,n+2)$, $n\geq 1$. It is therefore natural to ask whether the method of Kohnen and Skoruppa can be generalised to the orthogonal setting.
A first attempt at such a result was recently made by one of the authors in [psyroukis_orthogonal], although it differs in several fundamental respects from the present work. To start with, [psyroukis_orthogonal] only considers a single Dirichlet series, corresponding to a modular form $F$ of full level. Moreover, the condition that the algebraic subgroup $H(\xi)$ has class number one must be imposed in order to establish a clear-cut Euler product. These conditions substantially restrict the range of admissible orthogonal groups.
This paper serves as a generalisation of the work [psyroukis_orthogonal] for the case of signature $(2,4)$ in two crucial ways. First, we are able to remove the class number one assumption on the group $H(\xi)$ by exploiting the accidental isogeny between $\textup{U}(2,2)$ and $\textup{SO}(2,4)$. Moreover, we introduce–for the first time–the idea of considering multiple Dirichlet series, roughly corresponding to the number of classes in the genus of the definite group corresponding to $\phi$. This refinement is essential in order to be able to prove our result for number fields $K$ of any class number. We should mention here that most of the ideas discussed up until Section 6 can be generalised for arbitrary $n\geq 1$. We expect that a similar approach, considering multiple Dirichlet series, should be the appropriate method to obtain a clear-cut Euler product in the general case.
Finally, our result generalises another classical result, namely that of Gritsenko [gritsenko]. In particular, Gritsenko generalised the result of Kohnen and Skoruppa for the case of Hermitian modular forms of degree two over the imaginary quadratic field $K=\mathbb{Q}(i)$. Gritsenko’s approach, which relies on factorisation methods in parabolic Hecke rings, is very different from the techniques used in this paper. Indeed, our work is closer to the approach of Kohnen and Skoruppa [kohnen_skoruppa], in which the representation numbers of binary Hermitian forms play an analogue role to those of binary quadratic forms. To this, we should add that it is not difficult to see that Gritsenko’s method can be generalised to any number field $K$ of class number $1$. However, to the best of our knowledge, nothing is known, or at least published, for the case of a class number larger than one. In that sense, the present work may also be viewed as a generalisation of Gritsenko’s results to arbitrary number fields as in Theorem 12.6, due to the accidental isogeny $\textup{U}(2,2)$ and $\textup{SO}(2,4)$.